OpenStudy (anonymous):

write the equation on the line that passes through point (1, -4) and has a slope of 6

OpenStudy (anonymous):

(y-y1)=m(x-x1) where m is your slope. Then plug your values for (x1,y1) from above and solve

OpenStudy (anonymous):

can you do it cause im too lazy pleaseee

OpenStudy (fibonaccichick666):

^that is called point-slope form. It is very useful

OpenStudy (fibonaccichick666):

No. no. no. please see the code of conduct.

OpenStudy (anonymous):

i suck at math tbh but hey i tried! is it Y + 4 = 6 (x - 1)

OpenStudy (anonymous):

That's the first step...Now you need to simplify the equation

OpenStudy (anonymous):

Y + 6x - 10 ?

OpenStudy (anonymous):

Wait I meant Y = 6x - 10

OpenStudy (anonymous):

Great Job

OpenStudy (anonymous):

Yasss thanks broooo